# Ever-Evolving Memory by Blending and Refining the Past

### 03 Mar 2024 v1

@@ Abstract
For a human-like chatbot, constructing a long-term memory is crucial. A naive approach for making a memory could be simply listing the summarized dialogue. However, this can lead to problems when the speaker’s status change over time and contradictory information gets accumulated. It is important that the memory stays organized to lower the confusion for the response generator. In this paper, we propose a novel memory scheme for long-term conversation, CREEM. Unlike existing approaches that construct memory based solely on current sessions, our proposed model blending past memories during memory formation. Additionally, we introduce refining process to handle redundant or outdated information. This innovative approach seeks for overall improvement and coherence of chatbot responses by ensuring a more informed and dynamically evolving long-term memory.
Ever-Evolving Memory by Blending and Refining the Past
@@ 1 Introduction
Humans undergo continuous changes in emotions and states over time. For instance, one who was into classical music might turn into a passionate fan of heavy metal, or one who was suffering from severe illness could now be healthy enough to jog every morning. In order to create a chatbot that is sensitive to these changes, it is crucial that the changes are immediately reflected in its long-term memory Adiwardana et al. (2020). In this paper, we introduce a memory model that seamlessly blends and refines past memories. We define two problems to make an advanced long-term memory that resolves the issues of the existing memory frameworks:
@@
Figure 1: Motivating examples where our memory (Blend + Refine) is needed compared to not refining (Case 1) and removing all contradictories (Case 2).
@@
[P1] Only the current session is considered when creating new memory.
Long-term memory can be constructed by simply storing the summarized dialogue or persona Jang et al. (2023); Xu et al. (2021). This approach, focusing solely on the current session, results in one-dimensional information, leading to unorganized and disconnected memories. Consequently, the integration of past memories becomes overly dependent on the response-generating language model.
[P2] Memory contains outdated or contradictory information.
Prior works have explored expanding memory beyond direct dialogue information using commonsense knowledge or high-level insights from recent memory Majumder et al. (2020); Kim et al. (2023); Park et al. (2023). However, these methods risk introducing excessive redundancy or contradictions. Figure 1 illustrates Case 1, where outdated memory might be prioritized, neglecting the current context. Additionally, simply removing contradictions Bae et al. (2022) during the refining process may result in missing information, leading to awkward conversations, as depicted in Case 2 of Figure 1.
To address these issues, we propose a novel memory framework Contextualized Refinement based Ever-Evolving Memory (CREEM). Our goal is to make an ever-evolving memory that effectively integrates long-term memory, seamlessly connecting past and new information. CREEM continuously refines its memory every time there is a new information to remember. The main contributions of our work are as follows:
  • We demonstrate how the memory quality aligns with response generating performance.
@@ 2 Proposed Approach: CREEM
2.1 Tasks
Task 1. Blending Past Memory (P1)
We should take past memories into consideration while making a new memory to be added. Prior works that construct memory with the information extracted from a single session can be expressed as equation 1 where $\mathrm{m}_{s}$ and $\mathbf{D}^{s}$ are new memories and dialogue at current time $s$ respectively, and $f$ is the memory expanding function. On the contrary, our method utilizes past memories while constructing the memory itself as equation 2 where $\mathbf{M}_{[0,s-1]}$ is the set of past memories accumulated.
@@ 2.1 Tasks
$\mathrm{m}_{s}=f(\mathbf{D}^{s},\mathbf{M}_{[0,s-1]})$ (2)
Task 2. Refining Memory (P2)
When adding new memory, past memory should be refined to match the overall consistency. We refine the existing memory by judging whether they are redundant or updated.
In this section, we will explain the details of each step. The overall method flow is shown in Figure 2 and Algorithm 1.
2.2 Contextual Search
The initial step is searching past memories helpful in forming new ones. To search for contextually relevant memories, we formulate a query using ChatGPT-3.5-turbo  OpenAI (2023). The generated question serves to extract top-k information not only from past memories but also the ongoing dialogue, ensuring recent utterances are considered in the memory retrieval process.
Evolutionary Fundamental
Integration Consistency Sophistication Conciseness Relevance Concreteness Diversity
CC Dataset
$\text{SumMem}_{\text{MSC}}$ 3.16 4.11 2.59 3.72 3.85 3.10 3.24
$\textsc{ReBot}_{\text{CC}}$ 4.01 4.50 3.57 3.90 4.51 4.01 4.00
CREEM 4.05 4.64 3.86 3.95 4.61 4.00 4.03
MSC Dataset
$\text{SumMem}_{\text{MSC}}$ 3.44 3.95 3.02 3.25 4.09 3.82 3.72
$\textsc{ReBot}_{\text{CC}}$ 3.86 4.11 3.52 3.36 4.23 3.97 4.00
CREEM 4.02 4.19 3.88 3.38 4.40 4.01 4.07
Table 1: G-eval of memory based on 7 criteria with CC and MSC dataset.
2.3 Blend and Refine
The Blend process uses the top-k information previously retrieved. Given those information, $n$ memories are newly generated through integration and high-level inference.
Simultaneously, the Refine step evaluates the relationship between new and past memories. It categorizes the inference into three options: New if it contains entirely novel information, Redundant if it overlaps with past memories, and Updated if past information has changed. In the case of Redundant or Updated, LLM identifies which past memories to remove. After deletion, the new insights are added, and the model generates a response based on the current dialogue and the added insights.
1: Definition: $GQ=GenerateQuery$,
2:$R=Retrieval$,   $BR=Blend\&Refine$
3: $GR=GenerateResponse$
@@ 2.3 Blend and Refine
5: $query\leftarrow GQ\left(\mathbf{D}^{s}_{[1,t-1]}\right)$
@@
[…]
8:$Response\leftarrow GR\left(\mathbf{M}_{t}^{s},\mathbf{D}^{s}_{[1,t-1]}\right)$
@@
CREEM $\textsc{ReBot}_{\text{CC}}$ $\text{SumMem}_{\text{MSC}}$
What present + past present present
When per turn per session per turn
Who each speaker no distinction each speaker
Model ChatGPT (two-shot) distilled T5-base supervised encoder-decoder
Fine-tuned $\times$ $\checkmark$ $\checkmark$
Table 2: Comparison of memory constructing methods.
@@ 3.1 Datasets
We evaluate on two long-term chat datasets. Conversation Chronicles (CCJang et al. (2023) is a dataset generated by ChatGPT given a relationship, time interval and the storyline employ the narratives from SODA Kim et al. (2022). Multi-Session Chat (MSCXu et al. (2021) is a dataset created by human annotators through expanding dialogues from Persona-Chat Zhang et al. (2018), given a set of persona of previous sessions..
We focus on using these datasets mainly as a source of rich topics and conversation settings for our experiment, rather than absolute benchmarks. In order to validate the effect of a long-term memory, we experiment for the final sessions which has accumulated sufficient memory. For testing, we use 5 utterances per turn across 50 episodes from each datasets, yielding approximately 250 test data each.
@@ 3.2 Models and Settings
We compare our memory approach with two other methods introduced with the datasets. For the CC dataset, $\textsc{ReBoT}_{\text{CC}}$ summarizes each session’s dialogue with a distilled T5-base model Raffel et al. (2020), using these summaries as long-term memory for both speakers. In the case of the MSC dataset, they employ its summarization model, which we will call as $\text{SumMem}_{\text{MSC}}$, to extract persona for each speaker’s utterance. Table 2 provides a detailed comparison.
We set both baseline methods to receive the entire memory as input for response generation, following the original $\textsc{ReBoT}_{\text{CC}}$ approach. For CREEM, Contriever Izacard et al. (2021) is used as the retriever to extract 5 information pieces, and the Blend process generates 2 new memories. For fairness, we use ChatGPT as the response generation model for all methods.
3.3 Results and Discussion
@@ 3.3.1 Memory Evaluation
First of all, we should define what is a good memory? A good memory should consistently offer relevant information aligned with past conversations. Similar to human memory, it should evolve, keeping track of changes and resolving conflicts between old and updated information.
We establish 4 criteria for evaluating the qualities as a next-generation memory, under the evolutionary aspect: accurate connection and reflection of changes across sessions in details (Integration); absence of contradictory or conflicting information (Consistency); advancement beyond one-dimensional, direct information (Sophistication); well-organized, avoiding unnecessary repetition or extraneous details (Conciseness).
We set additional 3 criteria to assess the fundamental aspects of a memory: close relation to the speaker (Relevance); specific and precise information (Concreteness); variety of topics, perspectives, or insights (Diversity).
We conduct G-Eval Liu et al. (2023) customized to evaluate the memory given each of the criteria, instruction steps, and the conversation of all sessions.
Table 1 displays scores for three models across datasets. CREEM outperforms in all four evolutionary aspects, indicating its superior ability to integrate, maintain consistency, achieve sophistication, and ensure conciseness. The Blend process appears to play a crucial role in integrating information across sessions, creating sophisticated memories by expanding the memory. Additionally, the Refine process contributes to maintaining a consistent and well-organized memory by continually removing redundant or outdated information. Even in fundamental aspects, CREEM excelled in all criteria except for a slight difference in Concreteness in the CC dataset.
CREEM vs. $\textsc{ReBoT}_{\text{CC}}$ $\text{SumMem}_{\text{MSC}}$
CC Dataset
   Consistency 61.29 % 74.52 %
   memorability 48.06 % 70.97 %
MSC Dataset
   Consistency 63.85 % 67.57 %
   memorability 61.15 % 61.15 %
@@
CREEM
Original vs. ()(-)( - ) refine ()(-)( - ) pre-rg Blend
CC Dataset
   Consistency 64.41 % 53.55 %
   memorability 59.66 % 47.74 %
MSC Dataset
   Consistency 64.32 % 63.77 %
   memorability 63.40 % 58.70 %
Table 4: Results of ablation study. We report CREEM’s winning rate.
3.3.2 Response Generation Evaluation
@@
We conduct a pair-wise evaluation with ChatGPT based on 2 criteria that can most effectively show the role of the memory: maintains consistency without contradictions from past sessions (Consistency); effectively recalls past events correctly by retaining information from previous sessions (Memorability).
Table 3 shows the winning rate of CREEM. With CREEM’s memory, LLM creates better response in terms of consistency. $\textsc{ReBoT}_{\text{CC}}$ seems to slightly excel in memorability within it’s own dataset. This is possible considering that the CC dataset itself is made with ChatGPT and $\textsc{ReBoT}_{\text{CC}}$ uses fine-tuned memory model.
3.3.3 Ablation Study

### 07 Apr 2024 v2

@@ Abstract
For a human-like chatbot, constructing a long-term memory is crucial. However, current large language models often lack this capability, leading to instances of missing important user information or redundantly asking for the same information, thereby diminishing conversation quality. To effectively construct memory, it is crucial to seamlessly connect past and present information, while also possessing the ability to forget obstructive information. To address these challenges, we propose CREEM, a novel memory system for long-term conversation. Improving upon existing approaches that construct memory based solely on current sessions, CREEM blends past memories during memory formation. Additionally, we introduce a refining process to handle redundant or outdated information. Unlike traditional paradigms, we view responding and memory construction as inseparable tasks. The blending process, which creates new memories, also serves as a reasoning step for response generation by informing the connection between past and present. Through evaluation, we demonstrate that CREEM enhances both memory and response qualities in multi-session personalized dialogues.
@@ 1 Introduction
To enhance the conversational experience with human-like chatbots, chatbots should possess long-term memory and utilize it to engage users in conversations. While today’s large language models (LLMs) have facilitated interactions with humans, they often lack long-term memory. This limitation leads to instances of forgetting important user information or redundantly exchanging information, ultimately diminishing the quality of conversation (Zhong et al., 2024). To bridge this gap and provide a more authentic conversational experience, constructing and integrating long-term memory into chatbot is imperative.
@@
In long-term conversations, the user’s situation and preference can continue to change as the number of sessions increases. For instance, one who was into classical music might turn into a passionate fan of heavy metal, or one who was suffering from severe illness could now be healthy enough to go jogging every morning. In order to create a chatbot that is sensitive to these changes, it is crucial that the changes are immediately reflected in its long-term memory (Adiwardana et al., 2020).
When constructing a memory system, the approach of simply storing the summarized dialogue or extracting persona may seem intuitive (Jang et al., 2023; Xu et al., 2021). However, these approaches do not effectively reflect changes in the conversation. They focus solely on the current session when extracting information to save, resulting in a listing of isolated information without capturing changes between sessions. Furthermore, as memory accumulates without a refining process, outdated or contradictory information may be intermixed (Bae et al., 2022). This hinders the retrieval of correct information, potentially leading to inappropriate responses as shown in Case 1 of Figure  1. Therefore, any memories that become useless or obstructive should be deliberately forgotten. While the refinement of memory is essential, research in this area remains insufficiently explored. To address both aspects effectively, we focus on the following question: How can we store past and new information without causing confusion or interference between them?
To this end, we propose a novel memory framework called Contextualized Refinement based Ever-Evolving Memory, CREEM. CREEM blends past and present information to generate insights, which are then added as new memories. Subsequently, refining operations are employed to forget outdated or redundant past memories. In contrast to traditional paradigms, we do not view response generation and memory construction as separate tasks. Our blending process, where new insights are created, serves to link the past and present. The insights inferred in this process are then infused to the response generation model, assisting in properly integrating the updated information. Therefore, the blending process serves as both a step for memory construction and reasoning for response generation.
To demonstrate the effectiveness of CREEM, we begin by conducting memory evaluation using G-Eval along with introducing new criteria. Additionally, we closely examine the impact of refining by assessing the level of contradiction using a natural language inference (NLI) model, and through questions that address changes in content across sessions. For response generation evaluation, we conduct both reference-based and reference-free evaluations. We observe that even without fine-tuning, our model displayed strong performance in memory augmented response generation.
In summary, our main contributions are as follows:
  • We enhance memory adapted responding ability by infusing the blended insight to the response generation model.
  • We demonstrate how the memory quality aligns with the response quality.
@@
Figure 1: Motivating examples where our memory (Blend + Refine) is needed compared to not refining (Case 1) and deleting all contradictories (Case 2).
2 Related Work
2.1 Long-term Memory Systems
The long-term memory system has been an ongoing subject of research. Xu et al. (2021) extracts persona by summarizing the utterances for each speakers and use them as a memory for generating response in ongoing conversation. They introduce Multi-Session Chat (MSC), a dataset created by human annotators through expanding dialogues from Persona-Chat (Zhang et al., 2018) In Jang et al. (2023), memory is constructed by summarizing the each session in order to remember the content of the previous session and have a conversation in the next session. They introduce a new dataset, Conversational Chronicles (CC), which is generated by ChatGPT given a relationship, time interval, and a storyline employing the narratives from SODA (Kim et al., 2022). These methods of building memory by summarizing conversations creates one-dimensional information, which may prevent high-level conversation or conversation appropriate to the situation. Meanwhile, our work creates new memories by integrating relevant past memories, rather than simply summarizing the current conversation. Research on more advanced versions of memory systems has also been conducted. Majumder et al. (2020) expands the user’s persona with the commonsense knowledge graph (Kim et al., 2023). In Park et al. (2023), its generative agent conducts high-level reflections based on observations of the world, and then add them to the memory. Motivated by this approach, we use GPT-3.5-turbo (OpenAI, 2023) to extract insights from the current dialogue to get high-level information.
2.2 Forgetting in Long-term Memory
When constructing long term memory, as sessions become longer, unnecessary or conflicting information may accumulate. Therefore, there is a need for a function in memory to forget information accordingly. Bae et al. (2022) propose a long-term memory management mechanism with memory refining operations. The operations remove or add memory by determining the relationship between past memory and the summary created from the current session. Among them, the ’delete’ operation deletes both memory when they are no longer needed to be remembered and the summary contains any content that is no longer true. Case 2 in Figure 1 shows how the deleting operation can affect the memory capability. We implement its mechanism to compare with our method through further evaluations. Zhong et al. (2023) introduce a memory updating mechanism, mirroring the Ebbinghaus Forgetting Curve theory. It selectively forgets and strengthens memories based on time elapsed and importance. However it does not handle contradiction in memory.
@@ 3 Proposed Approach: CREEM
In this section, we explain the details of each step that is executed every turn from constructing memory to generating responses. The overall method flow is shown in Figure 2 and Algorithm 1.
3.1 Contextual Search
The initial step involves contextual searching to gather a robust pool of past memories, which serve as a valuable source for blending when forming new ones. Typically, a dialogue context is utilized as a query to retrieve pertinent memories. We take a step further by generating a question that aligns with the current dialogue context and using it as an additional query. This approach enables us to retrieve a broader range of memories that are implicitly related, extending beyond those directly resembling the ongoing dialogue. We utilize GPT-3.5-turbo (OpenAI, 2023)(for short ChatGPT) for this process.
3.2 Blend and Refine
Blending.
When creating a new memory, past memories are taken into consideration to ensure coherence and relevance. The Blend process uses the top-k information that were previously retrieved to generate $n$ memories through integration and high-level inference. Prior works that construct memory with the information extracted from a single session can be expressed as equation 1 where $\mathrm{m}_{s}$ and $\mathbf{D}^{s}$ are new memories and dialogue at current time $s$ respectively, and $f$ is the memory expanding function. On the contrary, our method utilizes past memories while constructing the memory itself as equation 2 where $\mathbf{M}_{[1,s-1]}$ is the set of past memories accumulated.
@@ 3.2 Blend and Refine
$\mathrm{m}_{s}=f(\mathbf{D}^{s},\mathbf{M}_{[1,s-1]})$ (2)
Refining.
When adding new memory, past memory should be refined to match the overall consistency. During the Refine step, the relationship between new and past memories is evaluated and categorized into three types: New if it introduces entirely novel information, Redundant if it overlaps with existing memories, and Updated if past information has been updated. Subsequently, based on the type of insight, appropriate actions are taken regarding past information. In the case of Redundant or Updated, LLM identifies which past memories to remove. After deletion, the new insights are added to the memory stream.
3.3 Insight Infused Response Generation
To generate responses, we provide the ongoing dialogue and the retrieved past memories that have undergone the refining process, along with newly created insights. The generated insights reflect changes or supplemented information based on the current utterance, thereby encapsulating connections between past and present. Through those insights, we enable the response generation model to comprehend the connections that embody changes between past and present. This ensures that responses accurately reflect the changes without any inconsistencies.
[…]
2: $s=\textit{session number}$, $t=\textit{turn number}$,
[…]
@@ 3.3 Insight Infused Response Generation
5: $query\leftarrow GenerateQuery\left(\mathbf{D}^{s}_{[1,t-1]}\right)$
@@
7: $\mathbf{M}_{t}^{s},type,\mathbf{M}_{remove}\leftarrow Blend\&Refine(\mathbf{[\textit{m'},...]},\mathbf{D}^{s}_{[1,t-1]})$
8:$Response\leftarrow GenerateResponse\left(\mathbf{M}_{t}^{s},\mathbf{[\textit{m'},...]}_{refined},\mathbf{D}^{s}_{[1,t-1]}\right)$
@@ 4.1 Datasets
We evaluate on two long-term chat datasets, Multi-Session Chat(MSC) (Xu et al., 2021) and Conversation Chronicles (CC) (Jang et al., 2023). While MSC is annotated by humans, CC is generated by ChatGPT. Both have 5 sessions of dialogue in each episode with various time intervals and persona. For testing, we use about 5 utterances per session across 50 episodes from each datasets, yielding approximately 250 sessions and 1250 utterances for each dataset.
$\text{SumMem}_{\text{MSC}}$ $\textsc{ReBot}_{\text{CC}}$ $\text{GPT}_{\text{CareCall}}$ CREEM
What present present present present + past
When per turn per session per session per turn
Model supervised encoder-decoder distilled T5-base ChatGPT ChatGPT
Fine-tuned $\checkmark$ $\checkmark$ $\times$ $\times$
Table 1: Comparison of memory constructing methods.
@@ 4.2 Models and Settings
We compare our memory approach with the methods introduced with the datasets and another method reflecting the refining process of Bae et al. (2022).
$\text{SumMem}_{\text{MSC}}$
For the MSC dataset, they employ its own encoder-decoder abstractive summarizer, which we will call as $\text{SumMem}_{\text{MSC}}$, to extract persona for each speaker’s utterance.
$\textsc{ReBoT}_{\text{CC}}$
For the CC dataset, $\textsc{ReBoT}_{\text{CC}}$ summarizes each session’s dialogue with a distilled T5-base model (Raffel et al., 2020), using these summaries as long-term memory for both speakers. The T5-base model is trained with the summaries of 80K sessions generated by ChatGPT.
We set the two baseline methods to receive the entire memory as input for response generation, following the original $\textsc{ReBoT}_{\text{CC}}$ approach.
$\text{GPT}_{\text{CareCall}}$
To evaluate our refining method, we experiment with the memory management mechanism proposed in Bae et al. (2022). Since there is no publicly available model, we implemented each operation using ChatGPT which we call $\text{GPT}_{\text{CareCall}}$. We implemented the methods using ChatGPT and Contriever to replicate the same model settings with our method for a fair comparison in the experiments. $\text{GPT}_{\text{CareCall}}$ performs pairwise processing of the existing memory and the summary generated at the end of the session, one by one; when memory already contains information of summary (PASS); when memory and summary are inconsistent, or when summary contains information of memory (REPLACE); when memory and summary are not related (APPEND); when memory is no longer true and summary is no longer needed to be remembered (DELETE).
CREEM
For CREEM, Contriever (Izacard et al., 2021) is used as the retriever to extract 5 information pieces each for question query and dialogue context query. Before each turn to respond, we instruct ChatGPT to generate 2 new insights to be added to the memory. We set the temperature parameter to 0.7 in order to create diverse insights. We provide detailed prompts in Appendix A for reproducibility.
Table 1 provides a detailed comparison. For fair comparison regarding the memory systems, we use ChatGPT as the response generation model for all methods.
4.3 Evaluation Metrics
@@ 4.3.1 Memory Evaluation
Before assessing memory performance, it is essential to establish the criteria for what defines an effective memory. A good memory should consistently offer useful information aligned with past conversations. Similar to human memory, it should evolve, keeping track of changes and resolving conflicts between old and updated information. We establish three criteria for evaluating the qualities as a next-generation memory, under the evolutionary aspects:
  • Integration: accurate connection and reflection of changes across sessions in details
  • Consistency: absence of contradictory or conflicting information
  • Sophistication: advancement beyond one-dimensional, direct information
We set additional three criteria to assess the fundamental aspects of a memory:
  • Relevance: closely related to the speakers
  • Concreteness: includes specific and precise information
  • Diversity: variety of topics, perspectives, or insights
We conduct G-Eval (Liu et al., 2023) customized to evaluate the memory given each of the criteria, instruction steps, and the conversation of all sessions.
Dataset Model Evolutionary Fundamental
Integration Consistency Sophistication Relevance Concreteness Diversity
MSC $\text{SumMem}_{\text{MSC}}$ 3.44 3.95 3.02 4.09 3.82 3.72
$\textsc{ReBot}_{\text{CC}}$ 3.86 4.11 3.52 4.23 3.97 4.00
$\text{GPT}_{\text{CareCall}}$ 3.86 4.69 3.33 4.14 3.52 3.44
CREEM 4.13 4.72 3.72 4.46 4.19 4.22
CC $\text{SumMem}_{\text{MSC}}$ 3.16 4.11 2.59 3.85 3.10 3.24
$\textsc{ReBot}_{\text{CC}}$ 4.01 4.50 3.57 4.51 4.01 4.00
$\text{GPT}_{\text{CareCall}}$ 3.88 4.87 3.67 4.30 3.59 3.21
CREEM 4.17 4.94 3.83 4.61 4.19 4.03
Table 2: G-Eval of memory based on 6 criteria with CC and MSC dataset.
4.3.2 Contradiction Level Evaluation
To conduct a more comprehensive evaluation focusing on the contradictory aspect, we assess the contradiction level by computing the average probability of contradiction for all pairs within the memory. This evaluation is performed using RoBERTa (Liu et al., 2019), which is fine-tuned on the MultiNLI(MNLI) dataset (Williams et al., 2017).
4.3.3 Question-answering Evaluation
To conduct a more robust evaluation of the refining process, we performed a question-and-answering(QA) based evaluation for CREEM and $\text{GPT}_{\text{CareCall}}$.
Question Generation
There have been studies of evaluating abstractive summarization through QA-based evaluation (Scialom et al., 2019; Wang et al., 2020). We create questions specifically designed to assess the memory system’s ability to remember altered information throughout the session. Utilizing the Chain-of-Thoughts (Wei et al., 2022) and form-filling technique, we instruct ChatGPT to generated multiple-choice questions. Given the entire session dialogue, ChatGPT identifies the changed information denoted as [Before] and [After]. Then, it creats questions based on this captured information. This process ensures that shifts in the speaker’s status to be accurately reflected within the problem. Additionally, a ”No Answer” option is included for cases where the question cannot be answered based on its memory. Please refer to the Appendix B.2 for detailed prompts.
Answering
The quality of memory is evaluated based on how well an answering model utilized the given memory to answer questions, measured by the accuracy rate of its responses. We utilize ChatGPT as the answering model, providing each memory as the sole external information.
4.3.4 Response Generation Evaluation
It is necessary to verify whether the quality of memory aligns with the quality of responses. To evaluate response quality, we conduct reference evaluations using BLEU, ROUGE-L, and BERTScore (Zhang et al., 2019; Papineni et al., 2002; Lin, 2004). However, a high score in these metrics does not necessarily imply that the model generated a superior response. Given the potential diversity in responses even within the same context, deviation from the reference response should not be interpreted as an incorrect answer.
Hence, alongside the automated evaluations, we perform a reference-free pair-wise evaluation using ChatGPT. Since we are focusing on the memory’s impact rather than the response generation performance itself, we established 2 criteria that can most effectively show the role of the memory:
  • Consistency: maintains consistency without contradictions from past sessions
  • Memorability: effectively recalls past events correctly by retaining information from previous sessions
Dataset Model Session 2 3 4 5 MSC $\text{SumMem}_{\text{MSC}}$ 0.3475 0.3514 0.363 0.3658 $\textsc{ReBot}_{\text{CC}}$ 0.198 0.2065 0.2157 0.2228 $\text{GPT}_{\text{CareCall}}$ 0.3054 0.2653 0.2916 0.2678 CREEM 0.1493 0.1476 0.1539 0.1613 CC $\text{SumMem}_{\text{MSC}}$ 0.1476 0.1656 0.1743 0.1775 $\textsc{ReBot}_{\text{CC}}$ 0.1717 0.177 0.2054 0.2006 $\text{GPT}_{\text{CareCall}}$ 0.1176 0.1036 0.1583 0.1241 CREEM 0.11 0.1454 0.1872 0.1915 Table 3: Degree of Contradiction. [Uncaptioned image] Figure 3: QA Accuracy Rate.
5 Results
5.1 Main Results
@@
Table 2 displays scores of G-eval over memory quality. CREEM outperforms in all three evolutionary aspects, indicating its superior ability to integrate, maintain consistency, achieve sophistication. The Blend process appears to play a crucial role in integrating information across sessions, creating sophisticated memories by expanding the memory. Additionally, the Refine process contributes to maintaining a consistent and well-organized memory by continually removing redundant or outdated information. Even in fundamental aspects, CREEM excelled in all criteria.
CREEM’s refining operations are effective in maintaining memory consistency.
Table 3 presents the degree of contradiction within the memory predicted by the NLI model. Overall, CREEM exhibited a low level of contradiction. However, concerning the CC dataset, $\text{GPT}_{\text{CareCall}}$ showed even lower levels of contradiction. While CREEM merge contradictory information into one, $\text{GPT}_{\text{CareCall}}$ remove both or one of the contradictory information through its operations, resulting in a higher removal of contradictory data. This approach may lead to excessive information loss, as evidenced by QA evaluation results in Figure 3. CREEM shows higher accuracy rates in both datasets. The lower accuracy rates for $\text{GPT}_{\text{CareCall}}$ suggest that it may have missed significant information during the refinement process. These results show that CREEM’s Blend and Refine process effectively incorporates both past and future information, maintaining better memory retention while reducing contradiction.
Model MSC CC
BLEU ROUGE-L BERTScore BLEU ROUGE-L BERTScore
$\text{SumMem}_{\text{MSC}}$ 1.392 0.143 0.866 4.053 0.175 0.876
$\textsc{ReBot}_{\text{CC}}$ 1.416 0.148 0.863 3.861 0.195 0.877
$\text{GPT}_{\text{CareCall}}$ 1.485 0.139 0.862 3.987 0.171 0.873
CREEM 1.610 0.150 0.865 4.667 0.198 0.880
Table 4: Performance in response generation by BLEU, ROUGE-L, and BERTScore
Dataset Criteria CREEM vs.
$\textsc{ReBoT}_{\text{CC}}$ $\text{SumMem}_{\text{MSC}}$ $\text{GPT}_{\text{CareCall}}$- Question -- Refine -- Insight Infuse
MSC Consistency 81.52 % 60.09 % 76.45 % 62.67 % 72.88 % 72.37 %
Memorability 71.01 % 50.00 % 68.12 % 73.63 % 70.34 % 69.30 %
CC Consistency 76.45 % 69.12 % 81.29 % 66.21 % 59.26 % 67.53 %
Memorability 54.52 % 56.87 % 72.58 % 49.68 % 58.02 % 54.32 %
@@
Table 4 presents the reference-based evaluation of the quality of responses generated by each model across different datasets. Under the BERTScore metric, the MSC data shows slightly better results with its own fine-tuned model, $\text{SumMem}_{\text{MSC}}$. Aside from that, it shows that CREEM generates better responses even without fine-tuning. Table 5 shows the winning rate of CREEM. Overall, ChatGPT generates better answers when utilizing CREEM’s memory, compared to three other memories. Particularly in terms of consistency, ChatGPT receives better evaluations, indicating that the Refine process effectively generates consistent responses aligned with past conversations. We can conclude that the performance of the response generation is positively influenced by the quality of the memory.
5.2 Ablation Study
Contextual searching using generated questions is effective for response quality.
Table  5 shows the performance difference between CREEM and its variations, excluding question generation and only using the current dialogue as the query in the Contextual Search process. Except for the memorablity in CC data, the original version of including both question and dialogue as query tends to make better responses. This show that searching with the generated question retrieves a wider range of information, which can contribute to creating a better memory.